The head of a golf club might contain 250 grams of titanium. How many moles is this?




1
Expert's answer
2022-04-25T06:12:04-0400

n = m / M

M (Ti) = 47.9 g/mol

n (Ti) = 250 / 47.9 = 5.2 mol